OpenStudy (anonymous):

If you have a 45.0 V battery connected to a 10.0 ohms resistor and a 50.0 ohms resistor in series, what will be the total power in the circuit? A. 50.0 W B. 33.8 W C. 203 W D. 405 W

OpenStudy (michele_laino):

current is: I=V(R_1+R_2)= 45/(50+10)=...? since resistors #1 and #2 are connected in series

OpenStudy (anonymous):

45/60 = 0.75?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, the requested power is: P=45*0.75=...watts

OpenStudy (michele_laino):

namely: power P= V*I

OpenStudy (anonymous):

33.75! so our solution is choice B? 33.8 W ?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thanks!!:)

OpenStudy (michele_laino):

:)